### v1

# Elucidating the representation of images within an unconditional diffusion model denoiser

Abstract
Generative diffusion models learn probability densities over diverse image datasets by estimating the score with a neural network trained to remove noise. Despite their remarkable success in generating high-quality images, the internal mechanisms of the underlying score networks are not well understood. Here, we examine a UNet trained for denoising on the ImageNet dataset, to better understand its internal representation and computation of the score. We show that the middle block of the UNet decomposes individual images into sparse subsets of active channels, and that the vector of spatial averages of these channels can provide a nonlinear representation of the underlying clean images. We develop a novel algorithm for stochastic reconstruction of images from this representation and demonstrate that it recovers a sample from a set of images defined by a target image representation. We then study the properties of the representation and demonstrate that Euclidean distances in the latent space correspond to distances between conditional densities induced by representations as well as semantic similarities in the image space. Applying a clustering algorithm in the representation space yields groups of images that share both fine details (e.g., specialized features, textured regions, small objects), as well as global structure, but are only partially aligned with object identities. Thus, we show for the first time that a network trained solely on denoising contains a rich and accessible sparse representation of images.
1 Introduction
Generative diffusion methods provide a powerful framework for sampling from probability densities learned from complex high-dimensional data such as images [1, 2, 3]. The key to their success lies in deep neural networks trained to estimate the score (the gradient of the log of the noisy image distribution), which is achieved by optimization on a denoising task. Training and sampling algorithms for these models have been extensively studied [4], but the score estimation properties that enable their spectacular image generation capabilities are not understood. Since denoising generally relies on distinguishing signal from noise, these score networks must somehow learn to identify and isolate patterns and structures found in their training data, so as to preserve them while eliminating noise. Here, we open the “black box” of a score network to study this internal representation.
We examine a UNet [5] trained (unconditionally) for image denoising on the ImageNet dataset. Prior work [6, 7, 8, 9, 10] has shown that activations from such a model can be extracted to perform downstream tasks – such as classification and segmentation – with considerable success. We aim instead to understand what features arise in the internal space of the network. We find that the network computes a low dimensional sparse set of activations in the output channels of the middle (innermost) network block. This may be summarized with a vector comprised of the spatial averages of these channels, which provides a representation of the underlying clean image. We verify that this representation is stable when computed on an image contaminated with noise over a broad range of amplitudes, consistent with the denoising objective on which it is trained. We develop an algorithm for stochastic reconstruction, using a reverse diffusion algorithm conditioned on this representation vector. This decoding procedure recovers a sample from a set of images whose representation is the same as that of the target image. Visually, the reconstructed images are similar in terms of both local and global patterns, revealing both the commonality and diversity of attributes encoded in the representation.
This representation exhibits a number of intriguing properties. Roughly, the channels summarized by the representation vector fall into two categories: non-selective channels, which capture common features present in many images; and selective channels, which are specialized for patterns that occur in only a small subset of images. As a consequence, the representation vectors of images lie within a union of low-dimensional subspaces, each of which is spanned by many of the common channels and a small subset of the specialized channels. We demonstrate that distances in the representation space are meaningful: The Euclidean distance between two representations predicts the symmetrized KL divergence between the two conditional distributions induced by the representations. Images whose representation vectors are similar are also semantically similar. Finally, we partition the images in the ImageNet validation set by applying a clustering algorithm to their representation vectors. The emergent clusters capture the general visual appearance of the images, sharing both fine details as well as global structure, but are only partially aligned with object category labels. Thus, the denoising objective alone, without any external conditioning, engenders learning of high level features that carry semantic information, albeit not necessarily object identities.
2 Implicit sparse image representation
Diffusion models learn image densities from data using a network that is trained for denoising. Specifically, given a noisy image $x_{\sigma}=x+\sigma z$, with $z\in\mathcal{N}(0,{\rm Id})$ a sample of white Gaussian noise, one trains a network $s_{\theta}(x_{\sigma})$ to minimize the squared error:
$\ell(\theta)=\mathbb{E}_{x,\sigma,z}\|x-\hat{x}({x_{\sigma})}\|^{2}=\mathbb{E}_{x,\sigma,z}\|{s}_{\theta}(x_{\sigma})-\sigma z\|^{2}.$ (1)
The optimal solution is the conditional mean, which can be directly related to the score of the underlying noisy distributions using a relationship published in [11], but generally referred to as "Tweedie’s formula" [12, 13]:
$\hat{x}(x_{\sigma})=\mathbb{E}[x|x_{\sigma}]=x_{\sigma}+\sigma^{2}\nabla_{x_{\sigma}}\log p_{\sigma}(x_{\sigma}).$ (2)
Thus, the trained network provides an approximation of the family of score functions for all $\sigma$. Reverse diffusion methods draw samples through iterative partial application of the learned denoiser, thereby using this approximate score to ascend the probability landscape [1, 2, 14, 3].
2.1 Spatial averages of activations
We adopt a convolutional UNet architecture [5], commonly used in diffusion models. The model consists of three main components: a set of encoder blocks (the downsampling path), a middle block, and a set of decoder blocks (the upsampling path) (see Fig 9 and Appendix A for more details). Our goal is to examine the activations of these hidden layers to understand how the noisy image is transformed to produce the score. To reduce the dimensionality, we examine the simplest summary of channel activations, $a_{j}(x_{\sigma})$, consisting of the spatial averages of all channels at the output layer of each block, notated by $\bar{a}_{j}(x_{\sigma})\in\mathbb{R}^{d_{j}}$. The components of the vector $\bar{a}_{j}$, which are always nonnegative due to half-wave rectification (ReLU) nonlinearities, carry information about the features encoded by the corresponding channels. Because of the convolutional nature of the network, the activations within each channel indicate the presence or absence of a feature in different spatial locations. Thus if a feature associated with channel $i\in(0,d_{j})$ is present somewhere in the input image, $\bar{a}_{j}(x_{\sigma})[i]$ will be positive, otherwise zero.
It is worth noting that the structure of the UNet architecture imposes certain properties on $\bar{a}_{j}$. Idealized convolutions are translation equivariant, hence the spatial averages of their activations are translation invariant w.r.t to the extracted features and do not carry information about location. For ergodic processes, these averages are approximations of statistical moments (expected value of nonlinear functions), and such measurements have been used for representation of visual texture [15, 16, 17, 18]. However, translation equivariance is imperfect for UNets (and most other convolutional networks) since it is violated by zero-padded boundary handling [19, 20]. This is especially prominent in the deeper blocks of the encoder, where the boundary influence encroaches on a larger spatial portion of the channels. In these cases, $\bar{a}_{j}$ can carry information about the location of features in addition to their presence. This effect coincides with the growth of the receptive field (RF) in the deeper blocks. As a result, $\bar{a}_{j}$ is expected to carry location information for larger features.
2.2 Denoising and channel sparsity
One of the difficulties in studying image representation in diffusion models is that it is not obvious where to look. We aim to locate the layers whose feature vectors, $\bar{a}$’s, well represent the clean image underlying the noisy input image. Such feature vectors would be rich enough to capture patterns and regularities of the image behind distortions caused by the noise in the input. Unlike Variational AutoEncoders (VAEs) [21], for which the encoder, bottleneck and decoder are well-defined and are assigned distinct roles through a dual objective function, the components of the UNet denoiser are only defined architecturally, and the optimization operates on a single end-to-end denoising objective. As a result it is not clear where the denoising takes place.
Some insight arises from considering the operation of denoisers designed in the pre-DNN era. Traditionally, image denoisers operate by transforming the noisy image to a latent space in which the true image is concentrated (sparse), and the noise is distributed (dense). Then the noise is suppressed and the image preserved. Finally, the latent representation is transformed back to the image space. This basic description captures the Wiener filter (operating in the frequency domain), and more modern thresholding denoisers which are typically applied within a multi-scale wavelet decomposition [22]. In all cases, more concentrated signal representations lead to better separation of noise and signal, hence superior denoising.
In the era of deep learning, neural networks optimized for denoising far outperform traditional solutions. Assuming the same principles apply, we expect to observe an increase in sparsity of the representation in portions of the network responsible for removing noise. To search for this locus, we measure the sparsity of $\bar{a}$ at the input and output layers of the blocks of a UNet trained on ImageNet. Here, sparsity of $\bar{a}\in\mathbb{R}^{d}$ is quantified by the normalized participation ratio (PR), the squared ratio $L_{1}$ and $L_{2}$ norms:
$\text{PR}(\bar{a})=\frac{\|\bar{a}\|^{2}_{1}}{d\ \|\bar{a}\|^{2}_{2}}.$
where $d$ is the ambient dimensionality of the space, i.e. number of channels in the layer. $\text{PR}\in(0,1]$, and provides a soft measure of dimensionality: small values reflect low dimensionality (sparsity), and large values indicate high dimensionality (density).
Figure 1: Channel sparsity of input and output layers of the image denoising UNet. Inset histograms show normalized participation ratios (PR), of the spatially averaged input channels (orange) and output channels (green) for individual blocks. Encoder blocks exhibit increases in PR (i.e., decreased sparsity), while the middle block and decoder blocks exhibit increases in sparsity. Blocks $\{E_{1},D_{1}\}$ are not included since they have only one input/output channel, receptively. This is evidence that encoder blocks extract features to isolate noise and signal, and middle block and decoder blocks preserve those channels containing signal while suppressing those containing noise. (Notation: encoder blocks $\{E_{k}\}$, middle block ($M$), decoder blocks $\{D_{k}\}$, downsampling ($d$), upsampling ($u$),“skip” connections. See Appendix A)
Figure 1 shows distributions of sparsity for input/output layers of blocks in the UNet denoiser. Both encoding blocks offer a substantial decrease in sparsity of $\bar{a}$, but we observe a stark increase of sparsity in the middle and decoding blocks. This suggests that removal of noise, through suppression of channels whose activations are primarily carrying noise, starts at the middle blocks, and then continues in the decoder blocks. The decoder block $j$, hence, denoises the features present in $a_{j}$ passed from encoder $j$, conditioned on the denoised features coming from below. This is evidence that the encoder transforms the noisy image to a hierarchical representation in preparation for denoising, by extracting image structures learned from the training set. Thus, the representation lends itself to investigation at the output layers of the middle and decoder blocks, where the features are exposed after removal of noise.
2.3 Robustness of representation to noise
Figure 2: Stability of $\bar{a}$ across noise levels, for different network blocks. Plots show cosine similarity of $\bar{a}(x_{\sigma_{1}})$ and $\bar{a}(x_{\sigma_{2}})$, for $\sigma_{1}=0.5$, as a function of $\sigma_{2}$. $\bar{a}$ is most stable in the middle block (M). Note that $\bar{a}$ collapses as $\sigma$ falls to zero, for which the denoiser should compute the identity function.
To elucidate the relationship between $x$ and $\bar{a}$, we need first to clarify the effects of noise. Noise introduces two types of variability in $\bar{a}$: 1) $\bar{a}$ depends on the noise amplitude, $\sigma$, and 2) at any given noise level, $\bar{a}$ will also depend on the particular noise realization in $x_{\sigma}$. In the context of reverse-diffusion sampling algorithms, the former translates to the evolution of representation with time. Not surprisingly, the variance in $\bar{a}$ grows with noise level. To remove this variability, we take the mean of $\bar{a}$ across noise realizations, $\mathbb{E}_{z}\left[\bar{a}(x+\sigma z)\right]$. Fig 2 shows the effect of noise level on $\bar{a}$. Variability due to noise level depends on the block depth, but interestingly, $\bar{a}_{j}$ is most stable in the middle block: increased noise level mostly increases the amplitude but not the direction of $\bar{a}_{j}$. The noise resilience of $\bar{a}$ at the output layer of the middle block makes it a good candidate for study. Thus, for the remainder of this paper we focus on the representation in this layer, which we notate as:
$\phi(x_{\sigma})=\mathbb{E}_{z}\left[\bar{a}_{4}(x+\sigma z)\right].$
3 Stochastic image reconstruction by conditional sampling
In this section, we propose an algorithm to “decode” $\phi$. As with any other representation, decoding $\phi$ is essential to understanding what it represents. Computing $\phi$ involves taking spatial average of the activations, that renders the representation translation invariance. Hence, $\phi$ represents not only the image it is computed from, but a whole set of images which have different activations with the same spatial average. In other words, $\phi$ is a many to one function. We aim to sample from the density embedded within the denoiser (or equivalently the score network) using a diffusion algorithm, while requiring that all samples share the same $\phi(x^{c})$ computed from a target image, $x^{c}$. This is a conditional sampling solutions for a non-linear inverse problem. The forward model, however, is not an auxiliary network or external measurement function, but the down-sampling (Encoding) path of the same score network. Since the algorithm samples images that are consistent with the network’s own representation of a target conditioner image, we view it as implementing a stochastic reconstruction from the representation.
The algorithm is obtained by augmenting a reverse diffusion algorithm with a soft projection onto the non-convex set in pixel space defined by $\phi$. To achieve this, the score step is alternated with an iterative projection step, where the sample is modified until its representation matches $\phi(x^{c})$, analogous to methods used to generate texture images from their measured statistics [17]. The matching is implemented by minimizing the Euclidean distance between the sample and target representations, $\|\phi(x)-\phi(x^{c})\|^{2}$, which is achieved by back-propagating the gradient of the loss through the first half of the network (i.e. $\bar{a}_{4}(x_{\sigma})$). The algorithm, hence, consists of two alternating steps described in Algorithms 2 and 1: At every time point in the synthesis, first the $\phi$ of the sample is matched to the target’s, via back propagating a gradient in the score network. Second the score-directed step pushes the sample closer to $p(x)$ by removing noise.
Algorithm 1 Stochastic reconstruction from the representation
score network $s$, conditioner image $x^{c}$, sigma schedule $(\sigma_{T},...,\sigma_{0})$, distribution mean $m$
Draw $x_{T}\sim\mathcal{N}(m,\sigma_{T}^{2}\mathrm{Id})$triangleright Initialize sample
for $t\in(T,...,1)$ do $\triangleright$ Alternate between matching and score steps
    Draw $z\sim\mathcal{N}(0,\mathrm{Id})$
    $x_{t}^{c}=x_{c}+\sigma_{t}z$
    $x_{t}\leftarrow$ Guidance step $(x_{t},x_{t}^{c})$triangleright Update $x_{t}$ to match representations using Algorithm 2
    Draw $z\sim\mathcal{N}(0,I)$
    $x_{t}\leftarrow x_{t}+s(x_{t})+\sigma_{t-1}z$triangleright Update $x_{t}$ using the score
end for
$x\leftarrow x_{{}_{1}}+s(x_{1})$
return $x$
Algorithm 2 Guidance step
score network $s$ to compute $\phi(x)$, sample $x_{t}$, conditioner image $x^{c}_{t}$, learning rate of optimizer $\eta$
$\phi_{x^{c}_{t}}=\phi(x^{c}_{t})\hskip 5.69046pt\text{and}\hskip 5.69046pt\phi_{x_{t}}=\phi(x_{t})$triangleright Compute representations
$\mathcal{L}(x_{t})=\|\phi(x_{t})-\phi_{x^{c}_{t}}\|^{2}$triangleright Compute distance between representations
while not converged do $\triangleright$ Minimize distance by backpropagating gradient through $e$
    $\nabla_{x_{t}}\mathcal{L}(x_{t})=(\phi_{x^{c}_{t}}-\phi(x_{t}))^{T}\nabla_{x_{t}}\phi(x_{t})$triangleright Compute gradient
    $x_{t}\leftarrow x_{t}-\eta\nabla_{x_{t}}\mathcal{L}(x_{t})$triangleright Update
end while
return $x_{t}$
This procedure can be described as a guided sampling algorithm [23], where the synthesis is guided by the network’s own representation. Matching the $\phi$’s changes the trajectory by forcing the sample to be within the set of images whose $\phi$ is equal to the conditioner’s. For a visualization of the effect matching on $x_{\sigma}$ in one step see Figures 11 and 12. The matching step does not change the noise level (Figure 13), so it is in a sense "orthogonal" to the score step. This means that matching iterations guide the sample in the domain of images that are high probability according to $p_{\sigma}(x_{\sigma})$ to reach the set of images defined by $\phi(x_{\sigma}^{c})$.
Figures 3 and 10 show samples generated by the algorithm. In each panel, the representation is obtained from an image selected from the validation set of ImageNet64 [24]. Then the stochastic reconstruction algorithm is used to generate $8$ images with the same $\phi$. The conditionally generated samples are not identical, but all are visually similar to the conditioner image $x^{c}$. Importantly, they share location-specific global structure as well as location-nonspecific details. Thus, the visual commonalities and diversities in the samples reveal what is and is not captured by the spatial averages of the feature vector in the middle block.
Figure 3: Samples from the stochastic reconstruction algorithm 1. Image at center of each panel is a target image drawn from the validation set of ImageNet64. The surrounding eight images are reconstruction samples, conditioned on the representation vector $\phi$ obtained from the target image. Samples in each panel are semantically similar. The diversity of samples varies for different target images, demonstrating that the sets defined by different $\phi$’s are drawn from neighborhoods of different sizes. More examples in Appendix B.
4 Properties of the representation
In the previous sections, we extracted a sparse representation from the output layer of the middle block and developed a stochastic algorithm to recover an image from it. In this section, we explore some properties of $\phi$. In particular, we examine the components of this vector, and their relationship to the content of the (clean) image $x$. Then, we study semantic similarities in pixel space in relation to Euclidean distances in the representation space.
4.1 Specialized versus common channels
For each channel, $i$, in the output layer of the middle block we quantify its (non-)selectivity using the participation ratio:
$\frac{\|c(i)\|_{1}^{2}}{n\ \|c(i)\|_{2}^{2}}$
where $c(i)=\left(\phi({x_{1}}_{\sigma})[i],...,\phi({x_{n}}_{\sigma})[i]\right)$ is the concatenation of the $i^{th}$ entry of $\phi$ for all images in the ImageNet64 validation set (approximately $50k$ images). This provides an estimate of the number of active channels, with smaller values indicating more selectivity. The threshold is set according to activation at zero noise level (Details in Appendix C). Figure 4 shows the distribution of selectivity for all $512$ channels in the middle block. The distribution is bimodal. Channels fall roughly into one of the two categories: selective channels that respond only to a small fraction of images in the set, and non-selective channels that respond to most images.
Figure 4: Channel selectivity. Left: Participation ratios for each channel over ImageNet. Distribution is bimodal, correponding to channels that are highly specialized (and infrequently active) on left, and commonly used on right. Right: The panels show the set of images that maximally activate each of four particular specialized channels, revealing selectivity for rectangular periodic lattices, a bird on a branch, full-frame cylindrical objects, and dog faces. Sorted by selectivity $i={87,143,187,287}$.
Specialized channels respond to specific features or patterns, and have negligible responses for most images in the dataset [25]. Figure 4 shows sets of images that maximally excite of four example specialized channels. Images within each group clearly share similar patterns. The observed sparsity of $\phi$ is a consequence of the existence of specialized channels. Since each image only contains a subset of patterns out of all the patterns present in the data distribution, only a handful of specialized channels are activated for each image. Figure 5 shows an example image along with the top three specialized channels it activates. Each channel is selective for a particular property of the image. Over the entire dataset, all channels are used, but different channels are used for different images.
Figure 5: Specialized channels capture visual attributes and composition of an image. Left: Example image that activates several specialized channels. Right: Each panel shows the set of images that maximally activate one of the specialized channels activated by the example image, corresponding to images of people, periodic texture patterns, and images with left-right reflective symmetry. All three elements are present in the example image.
The interpretation of the non-selective (“common”) channels is less straightforward. They are on average larger in magnitude, respond to a larger set of frequencies and orientations, and are the most stable w.r.t to noise level. These imply that they capture shared or common image features, such as brightness, global structure of a scene, etc.
4.2 Distances in representation space
An immediate consequence of sparsity and channel selectivity is that the $\phi$’s lie in a union of subspaces in $\mathbb{R}^{d}$, whose dimensionality lies approximately in the range $[0.2d,0.3d]$ (see Figure 1). How many of these subspaces are used? On one extreme, if the zero entries of $\phi$ coincide for all images $x$, this number could be as small as one. On the other extreme, it is bounded above by the number of possible $k$-dimensional subspaces, approximately $\sum_{k=100}^{150}\binom{500}{k}$, which far exceeds the size of any feasible dataset. The assumption behind this extreme case is that features corresponding to entries of $\phi$ are independent from one another, so that all combinations of features are equally likely. We know intuitively that this is not the case. Only certain combinations are likely and frequent in the training data distribution. This implies that collections of $\phi$’s should lie in common subspaces (Figure 6).
Figure 6: Union of subspaces. Left: Two sets of images whose $\phi$’s lie on two subspaces. Middle/Right: Three components of the $\phi$ vectors (out of the 512) for these images. The vertical axis corresponds to a common channel, while the other two correspond to specialized channels, each selective for only one image cluster. As a result, the $\phi$ vectors lie on a union of two-dimensional subspaces in the displayed three dimensional ambient space.
The network is trained solely with a denoising objective, without labels, augmentations, or regularization to induce any specific type of similarity or grouping (e.g membership in the same class). Therefore proximity in the representation space defines a fully unsupervised form of similarity in the image space. What do images whose $\phi$’s lie on the same subspace have in common? To examine this, we gathered images from the dataset with highest cosine similarity to a given target image in the representation space. Figure 7 shows two sets of those images. Images that are closest in the $\phi$ are visually and semantically similar, in stark contrast to images that are closest in the pixel space.
Figure 7: Nearby $\phi$’s correspond to images that are semantically similar. Left: Two sets of images, each showing a target image (upper left), and a set of images whose $\phi$’s are closest to that of the target image in terms of cosine similarity. For the butterfly image, proximity in latent space is aligned with the object identity. For the tray of food, proximity is aligned with images depicting collections of similar objects. Middle: Same as left, but showing images closest to the target in terms of cosine similarity in the pixel domain. Right Symmetrized KL divergence between the distributions of conditional samples, plotted against Euclidean distance in the representation, for randomly selected pairs of $\phi$ vectors.
The observation that Euclidean distances in representation space are related to perceptual similarities in the image space can be expressed more precisely. Each $\phi$ characterizes a conditional density, ${p}(x|\phi)$ from which Algorithm 1 draws samples. Distances between these conditional densities should govern the separation of $\phi$’s. We show that Euclidean distances between $\phi$’s are correlated with distances between conditional densities they induce. This is captured by a Euclidean embedding property, which ensures that the separation of $\phi$’s is related to a distance between the probability distributions ${p}(x|\phi)$, and hence that there exists $0<A\leq B$ with $B/A$ not too large, such that
$\forall x_{1},x_{2}\leavevmode\nobreak\ \leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ A\|\phi(x_{1})-\phi(x_{2})\|^{2}\leq d^{2}({p}_{1},{p}_{2})\leq B\|\phi(x_{1})-\phi(x_{2})\|^{2}.$ (3)
We establish a distance between two conditional densities as
$d^{2}({p}_{1},{p}_{2})=\int_{0}^{\infty}$ $\Big{(}\mathbb{E}_{{p}_{1}}\left[\|\nabla\log{p}_{1}-\nabla\log{p}_{2}\|^{2}\right]+\mathbb{E}_{{p}_{2}}\left[\|\nabla\log{p}_{1}-\nabla\log{p}_{2}\|^{2}\right]\Big{)}\,\sigma\,d\sigma.$ (4)
where ${p}_{1}={p}_{\sigma}(x_{\sigma}|\phi_{1})$ and ${p}_{2}={p}_{\sigma}(x_{\sigma}|\phi_{2})$. This distance is based on the difference in the expected score assigned to $x_{\sigma}$ by ${p}(x_{\sigma}|\phi_{1})$ versus ${p}(x_{\sigma}|\phi_{2})$, integrated across all noise levels. It provide a distance by symmetrizing the Kullback-Leibler divergence $\text{KL}(p\|q)$ between two distributions $p$ and $q$ proved in [26]:
[…]
The rightmost panel in Figure 7 confirms that, for a set of randomly selected pairs of images, Euclidean distances of their representations satisfy the Euclidean embedding inequality (3) with $\frac{B}{A}=2.9$.
4.3 Emergent clusters
Emergent clusters Human-labeled object classes Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Representation vectors, $\phi$, are concentrated within emergent clusters, less so with respect to object classes. Top Left: Euclidean distance between centroids of two clusters plotted against square root of their average variance along the line connecting the centroids, for 100 randomly selected cluster pairs. For all pairs, centroids are more than two standard deviations apart, and thus well-separated. t-SNE visualization of 10 randomly selected clusters. Top Right: Analogous plots for pairs of human-labeled classes, which are substantially more overlapping. Bottom: Each row contains images randomly selected from a cluster. Images within a cluster are visually similar and share global organization and semantic patterns, but they not necessarily from the same class. For example, dogs are clustered based on their pose, but not their breed. Thus, from a learned unconditional denoiser, a new notion of similarity emerges. More examples in Appendix C.
The fact that pairwise Euclidean distances in the $\phi$ space are meaningful implies that similar images cluster together within the subspace. To test this, we applied a K-Means clustering algorithm [27] to the $\phi$ vectors computed from the ImageNet dataset (See Appendix C). Figure 8 shows that within clusters the $\phi$’s are well concentrated around centroids, and the centroids are well separated: centroids of adjacent clusters are at least 2 standard deviations apart. The $\phi$’s of human labeled classes, by comparison, are less well separated. As a result, we expect images from different classes to appear within the same cluster. Examples are shown in Figure 8 for 9 clusters (see Appendix C for more).
Images within same clusters clearly share location-specific global scene structure, as well as location-nonspecific detailed elements (e.g., objects and their constituent parts, texture). However, in many cases, the similarity is not the object identity: a variety of circular objects appear in one cluster, athletes or dogs on a grass background in another, and a variety of sea creatures on a blue background in another. What connects images within the same cluster is "the gist of the scene" [28, 29, 30, 31, 32], which is only partially consistent with human labeled class memberships in ImageNet. Remarkably, these emergent clusters arise solely from learning the score, by minimizing a denoising objective. In the unconditional model, there is no external source of information to dictate proximity in the latent space. Importantly, this reveals that the type of patterns that are useful for computing the score are not object identities.
5 Discussion
Generative diffusion models have shown incredible success in learning and sampling from image densities. This feat is due to near minimum mean square error of denoising networks used in these algorithms. We hypothesized that these networks must construct an internal representation of image features that differentiate signal from noise, and we sought to elucidate that representation. We found that the vector of spatial averages of channels in the middle block of a trained UNet can provide a signature of this representation. Across a dataset of photographic images, the components of this vector are sparse – a relatively small subset of channels are active for each image – and thus the scores of complex image distributions have a low-dimensional structure, which is made explicit in these networks.
In addition to the sparsity of the representation vectors, we found that the geometry of the representation space is meaningful. Images with similar representations are visually similar and share location-specific global structure and location-nonspecific details. We developed a stochastic reconstruction algorithm, and showed that Euclidean distances in representation space are correlated with the symmetrized KL divergence of reconstruction distributions. Finally, we found that unsupervised clustering of the representation vectors yields well-separated groups of images that are semantically related, but only partially aligned with object identity. Thus, a network trained “bottom-up”, using only a denoising objective and no external conditioning, labelling, augmentation or regularization, can nevertheless learn a rich and accessible representation of image structure.
Appendix A Architecture, Dataset, Training
A.1 Architecture
We used the original UNet architecture [5] with a few modifications. The architecture is made up of 3 encoder blocks, 1 middle block, and 3 decoder blocks. The number of layers in each block is 2, 3, and 6 for the encoder, middle and decoder blocks. Each layer consists of $3\times 3$ convolutions with zero boundary handling, Layer normalization and ReLU. The number of channels is 64 in the first encoder block and then grows by a factor of two after each downsampling operation, and decreased by a factor of two after each upsampling operation. This results in 64, 128, 256 channels in the encoder blocks, 512 channels in the middle block, and 256, 128, 64 in the decoding blocks. The total number of parameters is $\sim 13m$.
The receptive field size at the end of the middle block is $84\times 84$ surpassing the spatial size of the input images in the ImagNet dataset. This is required for capturing large structures in the absence of attention blocks [19]. The depth of decoder blocks is increased in order to increase the expressivity of the denoising operations. This is again related to the size of the receptive field at the end of each decoder block with respect to the representation at the end of the middle block.
Figure 9: Modified UNet architecture used in our experiments.
A.2 Dataset
We used down-sampled version of ImageNet data set [24] , 3×,64×643$\times$,64$\times$ 643 × , 64 × 64, throughout the paper. The training set consists of $\sim 1.2m$ images and the validation set consists of $\sim 50k$ images, of objects, animals, scenes, etc. We did not use the class labels for training.
A.3 Training
The network was trained to minimize mean squared loss using Adam optimizer, with an initial learning rate of 0.001 with a decay of a factor of 2 every 100 epochs. Total number of epochs was set to 1000. The size of each batch was $1024$.
The standard deviation of noise was drawn randomly for each image from a $\frac{1}{\sqrt{\sigma}}$ distribution to emphasize the small noise levels, since we have observed that the it takes more epochs for the denoising MSE to plateau for small noise levels under uniform $\sigma$ distribution.
Appendix B Stochastic reconstruction algorithm
We build our sampling algorithm based on the algorithm in [33] which does not require the noise level and follows an adaptive step size schedule. Hence, this network does not take the noise variance as an input and is a blind denoiser. This setup has the advantage that simplifies the network, which is important for analysis of the internal layers of the model.
Figure 10: More example samples. The samples are visually similar to the target image at the center of each panel. Interestingly, the location of the large, long-ranging image patterns are persevered in the samples, while the fine structures and details are diverse in their location with respect to the target image. See caption of Figure 3.
Figure 11: A block diagram illustrating the matching step implementation and its effect. Top: One forward pass through the UNet without matching guidance. Bottom: One forward pass after matching representations in the middle block. The noise initialization, $x_{t}$, are identical in both passes, but the output of the network is vastly different for them, due to the effect of matching. Matching helps the large structures in the conditioner image appear in the output image. See more examples of this effect in Figure 12.
Figure 12: Effect of matching on the output of the denoiser. Top row shows 4 noise samples (akin to initialization in the sampling trajectory). Leftmost coloum shows 4 conditioner target images. In the middle, each row shows the result of matching the $\phi$ of the noise initialization with the target’s and then one pass through the denoiser (score step). The rightmost column shows the output of the same network without matching (only the score step).
Figure 13: Matching does not change the noise level $\sigma$. This plots shows the standard deviation of noise on the image before and after the matching step, for a collection of random test images, at different input noise levels. $\sigma$ after matching is measured by the norm of score divided by the square root of ambient dimensionality, ${\|s(x_{\sigma})\|}/{\sqrt{n}}$. This is an approximation of the strength of the remainder noise on the image by assuming a Gaussian prior.
Appendix C Properties of representation
C.1 Channel selectivity
Threshold.
To set the threshold, when computing the selectivity of channels, Figure 4, we compute the $\phi$’s at zero noise level for all dataset. In the ideal case, $\phi$ should be zero, but in practice some almost 1/3 of the channels are near-zero. This small amplitude can serve as a channel-wise threshold: For each channel, we set the threshold to be equal to maximum average activation for the channel at $\sigma=0$.
C.2 Clustering
We used python implementation of K-Means clustering algorithm, from sklearn package.
Clustering at different noise levels leads to similar results, consistent with stability of $\phi$ over noise levels shown in Figure 2. This is not true when noise is very small, since $\phi$ collapses to zero at small noise levels.
K-mean clustering algorithms are sensitive to initialization, so the assignments of images to clusters changes with initialization, but always leads to the same semantic grouping. Even initialization from the centroids of the pre-defined class labels results in similar clusters. This means that even when we give the algorithm a chance to cluster the images within the same class together, it pushes away from that and re-allocate the assignments such that the images are grouped together based on "the gist of the scene" as opposed to object identity.
Figure 14: Continused from Figure 8. Random images from different clusters are shown in each row. Different cars with the same orientation are clustered together, but similar car brands with different orientations are assigned to different clusters. (see rows 2 and 3).